# Tony Lindeberg

### v5

@@ Abstract
Keywords:
Receptive field Complex cell Orientation selectivity Elongation Affine covariance Vision
journal: arXiv preprint
1 Introduction
@@
To understand the functional properties of the visual system, it is essential to aim at bridging the gap between computational models on one side and neurophysiological measurements on the other side. Specifically, whenever possible, it is desirable to construct theoretically principled models, which could then lead to a deeper understanding of visual processing modules and also generate predictions for further biological experiments. One area where it indeed seems to be possible to bridge the gap between neurophysiological recordings and principled theory concerns the normative theory for visual receptive fields (Lindeberg 2021). This theory has been developed from principled assumptions regarding symmetry properties of an idealized vision system, and leads to a canonical family of linear receptive fields in terms of spatial derivatives of affine Gaussian kernels. Interestingly, the shapes of these idealized receptive fields do rather well correspond to the qualitative shapes of simple cells recorded by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008).
One of the components in this normative theory for visual receptive fields is the assumption that the family of receptive fields ought to be covariant under spatial affine transformations, so as to enable more robust processing of the image data as variations in the viewing conditions imply variabilities in the image data caused by natural image transformations (Lindeberg 2023; 2024b; 2025b). Specifically, if an observer views the same object from different distances and viewing directions, then the local image patterns will be deformed by the varying parameters of the perspective transformations, which to first order can be approximated by local affine transformations. In the area of computer vision, it has specifically been shown that the property of covariance under spatial affine transformations, referred to as affine covariance, enables more accurate estimation of surface orientation, compared to using only isotropic receptive fields that do not support affine covariance (Lindeberg and Gårding 1997).
@@
To implement affine covariance in a vision system corresponds to using receptive field shapes subject to different affine spatial transformations, and will thereby specifically imply that receptive fields ought to be present for different degrees of elongation. In a companion work (Lindeberg 2024a), we have indeed investigated the consistency of that affine covariant hypothesis with neurophysiological measurements of the resultant of orientation selectivity curves obtained by Goris et al. (2015). There, we showed that predictions of orientation selectivity curves and orientation selectivity histograms generated from idealized models of simple cells in terms of Gaussian derivatives are for idealized model of simple cells in reasonable agreement with biological orientation selectivity histograms, and thereby consistent with an expansion of the receptive field shapes over the degree of elongation in the primary visual cortex of higher mammals.
@@
The modelling of complex cells performed in (Lindeberg 2024a) was, however, based on very much simplified models in terms of pointwise non-linear combinations of responses of simple cells, and thus not involving any spatial integration of the non-linear combinations over extended regions in the image domain. Additionally, the modelling of complex cells in (Lindeberg 2024a) was based on input from simple cells up to only order two, and not involving simple cells up to order 4, which were shown to lead to better agreement between the predicted orientation selectivity histograms and the actual biological orientation selectivity histograms accumulated for simple cells by Goris et al. (2015). The subject of this paper is to investigate a set of extended models of complex cells, and to demonstrate that these models offer a potential to lead to better agreement with biological orientation selectivity histograms compared to previous work.
@@
In this way, we will thus specifically demonstrate that the resulting modelling of complex cells is also consistent with an expansion of the receptive field shapes over the degree of elongation, and in this way consistent with the wider hypothesis about affine covariant visual receptive fields, previously proposed in Lindeberg (2021; 2023; 2024b). Based on these results, to be presented below, we propose to include an explicit expansion over the degree of elongation of the receptive fields as an essential component when modelling the computational function of complex cells.
@@
While the specific models of complex cells to be considered in this treatment will be highly idealized, in terms of generalized energy models of the output of idealized simple cells, we propose that the conceptual extension of models of complex cells with regard to an expansion over the degree of elongation of the receptive fields ought to generalize to also other types of functional models of complex cells.
@@
More generally, we propose that the presented methodology made use of in this paper, of subjecting computational models of complex cells to similar probing tests as used for probing the orientation selectivity properties of biological complex cells, is important for understanding the theoretical properties of the computational mechanisms that are used for modelling biological complex cells. In this way, we will specifically demonstrate that complementing previous energy models of complex cells based on Gaussian derivatives with a spatial integration stage, as well as making use of spatial derivatives up to order 4 as opposed to previous use of spatial derivatives only up to order 2, has the potential of leading to better explanatory properties of orientation selectivity histograms of biological neurons, compared to the previous modelling approaches in (Lindeberg 2025a; 2024a).
@@
Thereby, the simulation work needed to reveal the qualitative properties of the computational models also becomes far lower compared to explicit simulations of networks of neural models, for which the result may also depending on the settings of the hyperparameters, and for which there may not be sufficient neurophysiological data available to tune the hyperparameters in a well-founded manner.
A general biological motivation for this study is that biological experiments often tend to reveal a variability of neurons in a variety of different respects. Connectivity analysis of the anatomy also tend to releval a convergence along axonal projections. By the presented theory for idealized models of complex cells, we provide an explanation for why there ought to be a variability in the degree of elongation of receptive fields, because of desirable affine covariance properties of an idealized vision system, which we here specifically apply to the modelling of complex cells.
@@
2 Methods
2.1 Related work
@@
Orientation selectivity properties of biological neurons have been studied by Watkins and Berkley (1974), Rose and Blakemore (1974), Schiller et al. (1976), Albright (1984), Ringach et al. (2002), Nauhaus et al. (2008), Scholl et al. (2013), Sadeh and Rotter (2014), Goris et al. (2015) and Sasaki et al. (2015). Biological mechanisms for achieving orientation selectivity have also been investigated by Somers et al. (1995), Sompolinsky and Shapley (1997), Carandini and Ringach (1997), Lampl et al. (2001), Ferster and Miller (2000), Shapley et al. (2003), Seriès et al. (2004), Hansel and van Vreeswijk (2012), Moldakarimov et al. (2014), Gonzalo Cogno and Mato (2015), Priebe (2016), Pattadkal et al. (2018), Nguyen and Freeman (2019), Merkt et al. (2019), Wei et al. (2022) and Wang et al. (2024). In this paper, our focus is, however, not on neural mechanisms, but on functional properties at a macroscopic level.
@@
Receptive field models of simple in terms of Gaussian derivatives have been formulated by Koenderink and van Doorn (1984; 1987; 1992), Young and his co-workers (1987; 2001; 2001) and Lindeberg 2013; 2021), and in terms of Gabor functions by Marcelja (1980), Jones and Palmer (1987a; 1987b) and Porat and Zeevi (1988). More extensive theoretical models based on Gaussian derivatives have also been expressed by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Hansen and Neumann (2008), Wallis and Georgeson (2009), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024).
@@
The taxonomy into simple and complex cells in the primary visual cortex was proposed in the pioneering work by Hubel and Wiesel (1959; 1962; 1968; 2005). More extensive analysis of properties of simple cells have then been presented by DeAngelis et al. (1995; 2004), Ringach (2002; 2004), Conway and Livingstone (2006), Johnson et al. (2008), Walker et al. (2019) and De and Horwitz (2021), and regarding complex cells by Movshon et al. (1978), Emerson et al. (1987), Martinez and Alonso (2001), Touryan et al. (2002; 2005), Rust et al. (2005), van Kleef et al. (2010), Goris et al. (2015), Li et al. (2015) and Almasi et al. (2020), as well as modelled computationally by Adelson and Bergen (1985), Heeger (1992), Serre and Riesenhuber (2004), Einhäuser et al. (2002), Kording et al. (2004), Merolla and Boahen (2004), Berkes and Wiscott (2005), Carandini (2006), Hansard and Horaud (2011), Franciosini et al. (2019), Lindeberg (2020), Lian et al. (2021), Oleskiw et al. (2024), Yedjour and Yedjour (2024) and Nguyen et al. (2024).
@@
There have been some neurophysiological studies reported that mention receptive fields with different aspect ratios (Tinsley et al. 2003, Xu et al. 2016). According to our knowledge, there have, however, not been any previously developed models of complex cells, that involve explicit expansions of the receptive field shapes over the degree of elongation of the receptive fields, or or more specifically that are able to match the neurophysiological orientation selectivity histograms accumulated by Goris et al. (2015).
2.2 Background theory
@@
In this section, we will describe basic properties of the idealized models for idealized models of receptive fields and their orientation selectivity properties, which we will then build upon and extend in Section 3.
2.2.1 Idealized models for spatial receptive fields
@@
For modelling simple and complex cells in the primary visual cortex, we will build upon the generalized Gaussian derivative model for visual receptive fields proposed in Lindeberg (2013; 2021) and further developed in Lindeberg (2023; 2024b; 2025a; 2025b).
$\kappa=1$ $\kappa=2$ $\kappa=4$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Distribution of affine Gaussian derivative receptive fields (for the preferred image orientation $\varphi=0$) over the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ between $1$ to $4$, from left to right. Here, the vertical scale parameter kept is constant $\sigma_{2}=4$, while the horizontal scale parameter is smaller $\sigma_{1}\leq\sigma_{2}$. (first row) First-order directional derivatives according to (1) for $m=1$. (second row) Second-order directional derivatives according to (1) for $m=2$. (third row) Third-order directional derivatives according to (1) for $m=3$. (fourth row) Fourth-order directional derivatives according to (1) for $m=4$. (Horizontal axes: image coordinate $x_{1}\in[-16,16]$. Vertical axes: image coordinate $x_{2}\in[-16,16]$.)
@@
$T_{\scriptsize\mbox{simple}}(x_{1},x_{2};\;\sigma_{\varphi},\varphi,\Sigma_{\varphi},m)\\ =T_{\varphi^{m},\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\sigma_{\varphi}^{m}\,\partial_{\varphi}^{m}\left(g(x_{1},x_{2};\;\Sigma_{\varphi})\right),$ (1)
@@
  • $g(x;\;\Sigma_{\varphi})$ is a 2-D affine Gaussian kernel with its shape determined by the covariance matrix $\Sigma_{\varphi}$
@@
In Lindeberg (2021), it was demonstrated that idealized receptive field models of this type do rather well model the qualitative shape of biological simple cells as obtained by neurophysiological measurements by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008).
Figure 1 shows examples of such receptive fields for different orders of spatial differentiation $m\in\{1,2,3,4\}$ and different values of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}\in\{1,2,4\}$ between here the vertical and the horizontal scale parameters $\sigma_{2}$ and $\sigma_{1}$, respectively, for the here preferred orientation $\varphi=0$ for the receptive fields. In addition to this illustrated variability over the degree of elongation, an idealized vision system should additionally comprise a variability over the preferred orientation $\varphi$ of the receptive fields, and possibly also over the overall size of the receptive fields, as further developed in Lindeberg (2024b).
@@
As a simplest possible extension to non-linear complex cells, an affine quasi-quadrature measure of the form (Lindeberg 2020 Equation (39))
@@
was studied in Lindeberg (2024a), where
@@
    $L_{\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\\ =T_{\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (4)
    $L_{\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\\ =T_{\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (5)
  • $C_{\varphi}>0$ is a weighting factor between first and second-order information, which based on a theoretical analysis in (Lindeberg 2018) is often set to $C=1/\sqrt{2}$.
This model is closely related to the energy model of complex cells proposed by Adelson and Bergen (1985) and Heeger (1992), as well as inspired by the fact that odd- and even-shaped receptive fields have been reported to occur in pairs (De Valois et al. 2000). The quasi quadrature serves as an approximation of a quadrature pair, as formulated based on a Hilbert transform (Bracewell 1999, pp. 267–272), although instead formulated in terms of affine Gaussian derivatives, which are then summed up in squares in to reduce the phase dependency, see Lindeberg (2020) for further details.
Concerning the validity of such an energy-based model of simple cells to model the computational function of complex cells, it is interesting to note that when Touryan et al. (2005) extracted the eigenvectors of second-order Wiener kernels to model the computational function of complex cells, the first two eigenvectors turned out the have spatial shapes that very well agree with the shapes of first- and second-order Gaussian derivatives; compare with Figure 5B in Touryan et al. (2005). Hence, even though an energy model of complex cells in terms of the output from a set of simple cells may not be able to span the full flexibility in terms of the possible computational functions of complex cells, an approximation of the computational function of a complex cell in terms of an energy model in such a way ought to be able to reveal essential properties of the computational functionalities of complex cells.
Figure 2: To measure the orientation selectivity properties of an idealized receptive field (here illustrated by the level curves of a first-order derivative of an affine Gaussian kernel in the horizontal direction with preferred orientation $\varphi=0$), we compute the response to a sine wave (here illustrated by a set of darker level lines corresponding to the spatial maxima and minima of the sine wave) with inclination angle $\theta$. (Horizontal axis: spatial coordinate $x_{1}$. Vertical axis: spatial coordinate $x_{2}$.) (Illustration reproduced from Lindeberg (2025a) OpenAccess.)
2.2.2 Orientation selectivity properties
In Lindeberg (2025a; 2024a), the orientation selectivity properties of these idealized models of simple and complex cells were investigated in detail, by computing the responses to sine wave functions of the form
@@ 2.2.2 Orientation selectivity properties
where $\theta\in[-\pi/2,\pi/2]$ denotes the inclination angle of the sine wave in relation to the preferred orientation $\varphi=0$ of the receptive field and $\beta$ denotes the phase, see Figure 2 for an illustration. Specifically, in Lindeberg (2025a; 2024a), it was shown that with
@@
3 Results
3.1 Generalized idealized models of complex cells

### v7

# Orientation selectivity properties for integrated affine quasi quadrature models of complex cells

@@ Abstract
Introduction
To understand the functional properties of the visual system, it is essential to aim at bridging the gap between computational models on one side and neurophysiological measurements on the other side. Specifically, whenever possible, it is desirable to construct theoretically principled models, which could then lead to a deeper understanding of visual processing modules and also generate predictions for further biological experiments. Regarding visual perception, it is in particular important to develop a good understanding of the receptive fields [^1] According to the pioneering work by Hubel and Wiesel [1, 2, 3, 4], a visual receptive field is defined as the region in visual space that can contribute to the response of a visual neuron. In this work, we follow a functional extension of that approach, by defining a visual receptive field as the computational function that computes the response of a visual neuron to the visual stimuli within the support region of the receptive field, that is over the region in the visual space that can evoke a response of the neuron. in the early layers of the visual hierarchy, which constitute the fundamental primitives that the higher layers in the visual hierarchy are based upon.
@@ Introduction
One area where it indeed seems to be possible to bridge the gap between neurophysiological recordings and principled theory concerns the normative theory for visual receptive fields in Lindeberg [5]. This theory has been developed from principled assumptions regarding symmetry properties of an idealized vision system, and leads to a canonical family of linear receptive fields in terms of spatial derivatives of affine Gaussian kernels. Interestingly, the shapes of these idealized receptive fields do rather well correspond to the qualitative shapes of simple cells recorded by DeAngelis et al. [6, 7], Conway and Livingstone [8] and Johnson et al. [9]; see Fig. 12–18 in Lindeberg [5] for comparisons between biological receptive fields and idealized models.
[…]
@@
To implement affine covariance in a vision system corresponds to using receptive field shapes subject to different affine spatial transformations, and will thereby specifically imply that receptive fields ought to be present for different degrees of elongation. [^3] For a visual neuron with a spatial receptive field, one can conceive that the receptive field has different amounts of spatial extent in different spatial orientations in image space. One way to characterize the amount of elongation of a receptive field is by measuring the spatial extent over all possible orientations and then forming the ratio between the extreme values of the spatial extent over all the image orientations. For receptive field models formulated in terms of affine Gaussian derivatives, as used in this work, the receptive field can specifically be decomposed into a smoothing stage with an affine Gaussian kernel followed by spatial derivative computations. For that affine Gaussian derivative model, the degree of elongation $\kappa$ can be defined as the square root of the ratio of the eigenvalues of spatial covariance matrix $\Sigma$ in the affine Gaussian kernel according to Eq. (2), as later formalized in Eq. (7). In a companion work (Lindeberg [14]), we have indeed investigated the consistency of that affine covariant hypothesis with neurophysiological measurements of the resultant of orientation selectivity curves obtained by Goris et al. [15]. There, we showed that predictions of orientation selectivity curves and orientation selectivity histograms generated from idealized models of simple cells in terms of Gaussian derivatives are for idealized model of simple cells in reasonable agreement with biological orientation selectivity histograms. Thereby, those results for simple cells are consistent with an expansion of the receptive field shapes over the degree of elongation in the primary visual cortex of higher mammals.
@@
The modelling of complex cells [^4] According to the taxonomy of neurons in the primary visual cortex by Hubel and Wiesel [1, 2, 3, 4], a visual neuron is said to be simple if it (i) has distinct excitatory and inhibitory subregions, (ii) obeys roughly linear summation properties and (iii) the excitatory and inhibitory regions balance each other in diffuse lighting. A visual neuron that does not obey these properties is said to be a complex cell. This characterization applies to any single individual visual neuron. In our comparisons to biological data to be performed later in this paper, we will then compare statistics over populations of biological complex cells, for relating gross properties of our proposed computational models to biological data. performed in Lindeberg [14] was, however, based on very much simplified models in terms of pointwise non-linear combinations of responses of simple cells, and thus not involving any spatial integration of the non-linear combinations over extended regions in the image domain. Additionally, the modelling of complex cells in Lindeberg [14] was based on input from simple cells up to only order two, and not involving simple cells up to order 4, which were shown to lead to better agreement between the predicted orientation selectivity histograms and the actual biological orientation selectivity histograms accumulated for simple cells by Goris et al. [15]. The subject of this paper is to investigate a set of extended models of complex cells, and to demonstrate that these models offer a potential to lead to better agreement with biological orientation selectivity histograms compared to previous work.
@@
In this way, we will thus specifically demonstrate that the resulting modelling of complex cells is also consistent with an expansion of the receptive field shapes over the degree of elongation, and in this way consistent with the wider hypothesis about affine covariant visual receptive fields, previously proposed in Lindeberg [5, 10, 14]. Based on these results, to be presented below, we propose to include an explicit expansion over the degree of elongation of the receptive fields as an essential component when modelling the computational function of complex cells.
@@
While the specific models of complex cells to be considered in this treatment will be highly idealized, in terms of generalized quadratic energy models of the output of idealized simple cells, we propose that the conceptual extension of models of complex cells with regard to an expansion over the degree of elongation of the receptive fields ought to generalize to also other types of functional models of complex cells.
@@
More generally, we propose that the presented methodology made use of in this paper, of subjecting computational models of complex cells to similar probing tests as used for probing the orientation selectivity properties of biological complex cells, is important for understanding the theoretical properties of the computational mechanisms that are used for modelling biological complex cells. In this way, we will specifically demonstrate that complementing previous energy models of complex cells based on Gaussian derivatives with a spatial integration stage, as well as making use of spatial derivatives up to order 4 as opposed to previous use of spatial derivatives only up to order 2, has the potential of leading to better explanatory properties of orientation selectivity histograms of biological neurons, compared to the previous modelling approaches in Lindeberg [16, 14].
@@
Thereby, the simulation work needed to reveal the qualitative properties of the computational models also becomes far lower compared to explicit simulations of networks of neural models, for which the result may also depend on the settings of the hyperparameters, and for which there may not be sufficient neurophysiological data available to tune the hyperparameters in a well-founded manner.
A general biological motivation for this study is that biological experiments often tend to reveal a variability of neurons in a variety of different respects. Connectivity analysis of the anatomy also tend to reveal a convergence along axonal projections. In this treatment, for idealized models of complex cells, we explain why there ought to be a variability in the degree of elongation, because of desirable affine covariance properties of an idealized vision system. Specifically, we demonstrate that this hypothesis is consistent with experimental results regarding orientation selectivity properties of biological neurons.
@@
Finally, we will use the results from the presented treatment (i) for stating a set of more general predictions in the section “Explicit predictions for further modelling of complex cells”, to be used as guide for further modelling of biological complex cells by mathematical models, and (ii) proposing a conceptual extension of the previous methodology for probing the orientation selectivity of biological neurons. The latter extension consists of instead of just recording the result for a single angular frequency for each image orientation instead performing a two-parameter variation over both the angular frequency and the image orientation, to be able to better reflect a variability in the degree of elongation between different individual biological neurons.
While the presented computational mechanisms are in the paper technically developed for the proposed family of affine quasi quadrature models of complex cells, we argue that corresponding additions of such computational mechanisms could be considered also for other types of theoretical and computational models of complex cells.
Methods
Related work
@@ Methods
Orientation selectivity properties of biological neurons have been studied by Watkins and Berkley [17], Rose and Blakemore [18], Schiller et al. [19], Albright [20], Ringach et al. [21], Nauhaus et al. [22], Scholl et al. [23], Sadeh and Rotter [24], Goris et al. [15] and Sasaki et al. [25]. Biological mechanisms for achieving orientation selectivity have also been investigated by Somers et al. [26], Sompolinsky and Shapley [27], Carandini and Ringach [28], Lampl et al. [29], Ferster and Miller [30], Shapley et al. [31], Seriès et al. [32], Hansel and van Vreeswijk [33], Moldakarimov et al. [34], Gonzalo Cogno and Mato [35], Priebe [36], Pattadkal et al. [37], Nguyen and Freeman [38], Merkt et al. [39], Wei et al. [40] and Wang et al. [41]. In this paper, our focus is, however, not on neural mechanisms, but on functional properties at a macroscopic level.
@@
Receptive field models of simple in terms of Gaussian derivatives have been formulated by Koenderink and van Doorn [42, 43, 44], Young and his co-workers [45, 46, 47] and Lindeberg [48, 5], and in terms of Gabor functions by Marcelja [49], Jones and Palmer [50, 51] and Porat and Zeevi [52]. More extensive theoretical models based on Gaussian derivatives have also been expressed by Lowe [53], May and Georgeson [54], Hesse and Georgeson [55], Georgeson et al. [56], Hansen and Neumann [57], Wallis and Georgeson [58], Wang and Spratling [59], Pei et al. [60], Ghodrati et al. [61], Kristensen and Sandberg [62], Abballe and Asari [63], Ruslim et al. [64] and Wendt and Faul [65].
@@
The taxonomy into simple and complex cells in the primary visual cortex was proposed in the pioneering work by Hubel and Wiesel [1, 2, 3, 4]. More extensive analysis of properties of simple cells have then been presented by DeAngelis et al. [6, 7], Ringach [66, 67], Conway and Livingstone [8], Johnson et al. [9], Walker et al. [68] and De and Horwitz [69], and regarding complex cells by Movshon et al. [70], Emerson et al. [71], Martinez and Alonso [72], Touryan et al. [73, 74], Rust et al. [75], van Kleef et al. [76], Goris et al. [15], Li et al. [77] and Almasi et al. [78], as well as modelled computationally by Adelson and Bergen [79], Heeger [80], Serre and Riesenhuber [81], Einhäuser et al. [82], Kording et al. [83], Merolla and Boahen [84], Berkes and Wiscott [85], Carandini [86], Hansard and Horaud [87], Franciosini et al. [88], Lindeberg [89], Lian et al. [90], Oleskiw et al. [91], Yedjour and Yedjour [92], Nguyen et al. [93] and Almasi et al. [94].
Notably, in relation to the generalized quadratic models of complex cells in V1 to be considered in this paper, Rowekamp and Sharpee [95] have found that quadratic computations strongly increase both the predictive power of their models of visual neurons in V1, V2 and V4 as well as their neural selectivity to natural stimuli.
@@
There have been some neurophysiological studies reported that mention receptive fields with different aspect ratios (Tinsley et al. [96], Xu et al. [97]). According to our knowledge, there have, however, not been any previously developed models of complex cells, that involve explicit expansions of the receptive field shapes over the degree of elongation of the receptive fields, or or more specifically that are able to match the neurophysiological orientation selectivity histograms accumulated by Goris et al. [15].
Background theory
@@
In this section, we will describe basic properties of the idealized models for idealized models of receptive fields and their orientation selectivity properties, which we will then build upon and extend in the section ’Results”.
Idealized models for spatial receptive fields
@@
For modelling simple and complex cells in the primary visual cortex, we will build upon the generalized Gaussian derivative model for visual receptive fields proposed in Lindeberg [48, 5] and further developed in Lindeberg [10, 16, 14, 12].
@@
$T_{\scriptsize\mbox{simple}}(x_{1},x_{2};\;\sigma_{\varphi},\varphi,\Sigma_{\varphi},m)=\\ =T_{\varphi^{m},\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\sigma_{\varphi}^{m}\,\partial_{\varphi}^{m}\left(g(x_{1},x_{2};\;\Sigma_{\varphi})\right),$ (1)
@@
  • $g(x;\;\Sigma_{\varphi})$ is a 2-D affine Gaussian kernel with its shape determined by the spatial covariance matrix $\Sigma_{\varphi}$
@@
In Lindeberg [5], it was demonstrated that idealized receptive field models of this type do rather well model the qualitative shape of biological simple cells as obtained by neurophysiological measurements by DeAngelis et al. [6, 7], Conway and Livingstone [8] and Johnson et al. [9].
Fig. 1 shows examples of such receptive fields for different orders of spatial differentiation $m\in\{1,2,3,4\}$ and different values of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}\in\{1,2,4\}$ between the here vertical and the horizontal scale parameters $\sigma_{2}$ and $\sigma_{1}$, respectively, for the here preferred orientation $\varphi=0$ for the receptive fields. In addition to this illustrated variability over the degree of elongation, an idealized vision system should additionally comprise a variability over the preferred orientation $\varphi$ of the receptive fields, and possibly also over the overall size of the receptive fields, as further developed in Lindeberg [14].
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Fig 1: Distribution of affine Gaussian derivative receptive fields (for the preferred image orientation $\varphi=0$) over the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ between 11 to $4$, from left to right. Here, the vertical scale parameter kept is constant $\sigma_{2}=4$, while the horizontal scale parameter is smaller $\sigma_{1}\leq\sigma_{2}$. (first row) First-order directional derivatives according to (1) for $m=1$. (second row) Second-order directional derivatives according to (1) for $m=2$. (third row) Third-order directional derivatives according to (1) for $m=3$. (fourth row) Fourth-order directional derivatives according to (1) for $m=4$. (Horizontal axes: image coordinate $x_{1}\in[-16,16]$. Vertical axes: image coordinate $x_{2}\in[-16,16]$.)
@@
As a simplest possible extension to non-linear complex cells, an affine quasi-quadrature measure of the form (Lindeberg [89] Eq. (39))
@@
was studied in Lindeberg [14], where
@@
    $L_{\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=T_{\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (4)
    $L_{\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=T_{\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (5)
  • $C_{\varphi}>0$ is a weighting factor between first and second-order information, which based on a theoretical analysis in Lindeberg [98] is often set to $C=1/\sqrt{2}$.
This model is closely related to the energy model of complex cells proposed by Adelson and Bergen [79] and Heeger [80], as well as inspired by the fact that odd- and even-shaped receptive fields have been reported to occur in pairs (De Valois et al. [99]). The quasi quadrature serves as an approximation of a quadrature pair, as formulated based on a Hilbert transform (Bracewell [100], pp. 267–272), although instead formulated in terms of affine Gaussian derivatives, which are then summed up in squares in to reduce the phase dependency; see Lindeberg [89] for further details.
Concerning the validity of such an energy-based model of simple cells to model the computational function of complex cells, it is interesting to note that when Touryan et al. [74] extracted the eigenvectors of second-order Wiener kernels to model the computational function of complex cells, the first two eigenvectors turned out the have spatial shapes that very well agree with the shapes of first- and second-order Gaussian derivatives; compare with Fig. 5B in Touryan et al. [74]. Hence, even though an energy model of complex cells in terms of the output from a set of simple cells may not be able to span the full flexibility in terms of the possible computational functions of biological complex cells, an approximation of the computational function of a complex cell in terms of an energy model in such a way ought to be able to reveal essential properties of the computational functionalities of complex cells.
Orientation selectivity properties
In Lindeberg [16, 14], the orientation selectivity properties of these idealized models of simple and complex cells were investigated in detail, by computing the responses to sine wave functions of the form
@@
where $\theta\in[-\pi/2,\pi/2]$ denotes the inclination angle of the sine wave in relation to the preferred orientation $\varphi=0$ of the receptive field and $\beta$ denotes the phase; see Fig. 2 for an illustration.
Fig 2: Configuration for probing the orientation selectivity of an idealized receptive field model. To measure the orientation selectivity properties of an idealized receptive field (here illustrated by the level curves of a first-order derivative of an affine Gaussian kernel in the horizontal direction with preferred orientation $\varphi=0$), we compute the response to a sine wave (here illustrated by a set of darker level lines corresponding to the spatial maxima and minima of the sine wave) with inclination angle $\theta$. (Horizontal axis: spatial coordinate $x_{1}$. Vertical axis: spatial coordinate $x_{2}$.) (Adapted from Lindeberg [16] OpenAccess.)
Specifically, in Lindeberg [16, 14], it was shown that with
@@
Results
Generalized idealized models of complex cells